# On the Isbell Problem

### v1

@@ 8. Squares and products
The characterization of $[\kappa]^{<\omega}\leq_{T}\mathcal{U}$ given by 2.15 opens the possibility of characterize the generic existence of ultrafilters non-Tukey above $[\kappa]^{<\omega}$, so our first question is:
Question 9.1.
How can the generic existence of ultrafilters non-Tukey above $[\kappa]^{<\omega}$ be characterized? Are there some familiar cardinal invariants involved? What about generic existence of Tukey top ultrafilters?
Regarding the generic existence of Tukey top ultrafilters, it is possible to prove that in the side by side Silver model, Tukey top ultrafilters exist generically, and actually, for each filter $\mathcal{F}$ of character less than $2^{\omega}$, it is possible to add $\omega_{1}$-many sets to $\mathcal{F}$ to make sure that any ultrafilter extending $\mathcal{F}$ is Tukey top.
Question 9.2.
Is it consistent that there is no nowhere dense ultrafilter but there is an ultrafilter non-Tukey above $[\omega_{1}]^{<\omega}$?
Question 9.3.
Is there a basically generated ultrafilter in Shelah’s model for no nowhere dense ultrafilters? What about only non-Tukey above $[\omega_{1}]^{<\omega}$?
Question 9.4.
@@
Let us define the high spectrum of the Tukey types of ultrafilters on $\omega$ as the set of cardinals $\kappa\geq\omega_{1}$ for which there is an ultrafilter such that $\mathcal{U}\equiv[\kappa]^{<\omega}$,

### v2

@@ 10. Squares and products
We begin by recalling the two questions of section 2:
Question 11.1.
Question 11.2.
Is there a non-Tukey top ultrafilter in Shelah’s model for no $\mathsf{nwd}$-ultrafilters? A basically generated ultrafilter?
The characterization of $[\kappa]^{<\omega}\leq_{T}\mathcal{U}$ given by 2.15 opens the possibility of characterizing the generic existence of ultrafilters non-Tukey above $[\kappa]^{<\omega}$, so our first question is:
Question 11.3.
How can the generic existence of ultrafilters non-Tukey above $[\kappa]^{<\omega}$ be characterized?
Let us recall that J. Brendle and J. Flašková have investigated the generic existence of $\mathscr{I}$-ultrafilters where $\mathscr{I}$ is an ideal on $\omega$, and introduced the cardinal invariant $\mathfrak{ge}(\mathscr{I})$ for a given ideal $\mathscr{I}$. They proved that generic existence of $\mathscr{I}$-ultrafilters is equivalent to the equality $\mathfrak{ge}(\mathscr{I})=\mathfrak{c}$. Under the current context, ultrafilters which are not Tukey above $[\kappa]^{<\omega}$ are characterized as exactly the same class as $\tau_{\kappa}$-ultrafilters, so the natural question is whether there is a similar characterization of generic existence in terms of some relation between appropriated cardinal invariants.
Question 11.4.
Is there a nice characterization of generic existence of Tukey top ultrafilters?
Regarding Question 11.4, it is possible to prove that in the side by side Silver model, Tukey top ultrafilters exist generically, and actually, for each filter $\mathcal{F}$ of character less than $2^{\omega}$, it is possible to add $\omega_{1}$-many sets to $\mathcal{F}$ to make sure that any ultrafilter extending $\mathcal{F}$ is Tukey top.
Question 11.5.
Is any basically generated ultrafilter Sacks indestructible?
Question 11.6.
@@
It may be possible that the following questions are still far away to be reached just from the current consistency tools and machinery. Despite this, they are still interesting. Let us define the high spectrum of the Tukey types of ultrafilters on $\omega$ as the set of cardinals $\kappa\geq\omega_{1}$ for which there is an ultrafilter such that $\mathcal{U}\equiv[\kappa]^{<\omega}$,
@@
12. Acknowledgments.
The first author was supported by the “Programme to support prospective human resources – post Ph.D. candidates” of the Czech Academy of Sciences, project L100192251, and by the Czech Academy of Sciences CAS (RVO 67985840), and also supported by Universidad Nacional Autónoma de México Postdoctoral Program (POSDOC). The research for this paper was started when the first author was a postdoc at Institute of Mathematics of Czech Academy of Sciences, and concluded after he became a postdoc in the Universidad Autónoma de México (UNAM), under the posdoctoral fellowship program “Elisa Acuña”. He is grateful to both institutions.